### v1

# AI-enhanced tuning of quantum dot Hamiltonians toward Majorana modes

Abstract
We propose a neural network-based model capable of learning the broad landscape of working regimes in quantum dot simulators, and using this knowledge to autotune these devices – based on transport measurements – toward obtaining Majorana modes in the structure. The model is trained in an unsupervised manner on synthetic data in the form of conductance maps, using a physics-informed loss that incorporates key properties of Majorana zero modes. We show that, with appropriate training, a deep vision-transformer network can efficiently memorize relation between Hamiltonian parameters and structures on conductance maps and use it to propose parameters update for a quantum dot chain that drive the system toward topological phase. Starting from a broad range of initial detunings in parameter space, a single update step is sufficient to generate nontrivial zero modes. Moreover, by enabling an iterative tuning procedure – where the system acquires updated conductance maps at each step – we demonstrate that the method can address a much larger region of the parameter space.
I Introduction
Majorana zero modes (MZMs) are emergent quasiparticles predicted to obey non-Abelian statistics, enabling topologically protected quantum information processing [Alicea2012, Beenakker2013, Sau2010, Oreg2010]. Their realization in hybrid superconductor-semiconductor systems requires precise control over local parameters such as chemical potentials, inter-dot couplings, and spin-orbit interactions [Leijnse2012, Lutchyn2018]. Small deviations from the so-called sweet-spot conditions of the Kitaev chain (KC) [Leumer2021] can destroy the topological gap and delocalize MZMs [Kitaev2001]. Moreover, it is not trivial to distinguish between MZMs and different nontopological zero-bias peaks  [Aguado2021, Sasaki2000, Pikulin2012, Kells2012]. Consequently, achieving robust Majorana states experimentally remains a major challenge, particularly in the presence of fabrication disorder and parameter noise.
To address this challenge, Fulga et al. [Fulga2013] proposed an adaptive tuning protocol for superconductor-proximitized chain of quantum dots (QDs) that simulates a KC generalized on non-uniform parameters distribution across dots, trying to demonstrate how gate-controlled voltages and superconducting phases can be adjusted to reach a topologically nontrivial regime. In this paper we revisit this problem trying to automate it using deep neural estimators. More recently, further works indicate that through appropriate tuning, it is possible to realize a KC with emerged MZMs using QDs coupled through elastic co-tunneling and crossed Andreev reflection  [Leijnse2012, Sau2012, Dvir2023, Bordin2025]. All these proposals utilize coupling with more readily available $s$-wave superconductors; however, there are also proposals employing $p$-wave superconductors [Mohseni2023, Ezawa2024].
At the same time, a different strategy for machine learning (ML)-assisted autotuning of QD-based quantum simulators [Borsoi2024, Mills2019, Shandilya2025] using transport measurements [Darulov2020, Lunczer2019, Zwolak2023, roux2025] is gaining significant interest. The use of transport measurements in the form of conductance maps [Donarini2024, Shandilya2025] to derive insights about the Hamiltonian system [Blonder1982, grepkowa2023, Wang2017, Gebhart2023, Thamm2024, Taylor2024disorderlearning, Taylor2025analysis] appears to be a promising path towards the automatic tuning of Hamiltonian parameters. In particular, inverting the measured conductance matrices to determine electrostatic potential disorder, using evolutionary optimization [Thamm2024], supervised neural networks (NNs) [Taylor2024disorderlearning, vandriel2024, Taylor2025analysis], or both [Taylor2025disordermitigation], is a natural first step that can allow precise tuning of Hamiltonian parameters towards MZMs. Although these methods already allow for successful mitigation of disorder in nanowires [Taylor2025disordermitigation, Thamm2024], they rely on indirect, non-differentiable cost functions and heuristic evolutionary searches, rather than learning the broader behavior of the underlying physical system.
Figure 1: Scheme of the QDs-based KC-simulator autotuning system (PINNAT). Black (blue) arrows show the system training (inference) path. Vision transformer model is fed with conductance maps to predict parameter corrections, that should decrease $\mathcal{M}$-based loss function, simultaneously increasing the probability of MZMs emergence.
In contrast, here we introduce an unsupervised physics-informed [Raissi2019, kovachki2023neural] NN-based auto-tuning (PINNAT) model with Majorana physics embedded in the loss function, and show that by proper training the model can memorize various structures on conductance maps which enables to build an efficient quantum simulator autotuning system. Our AI-enhanced adaptive tuning framework directly extends QDs simulator by leveraging vision transformer (ViT) [dosovitskiy2021] NN architecture. The scheme of the proposed approach is presented in Fig. 1 with information flow during the model training (black arrows) and evaluation (inference, blue arrows).
II Model
Following a lattice model from Ref. [Fulga2013], the effective superconductor-proximitized Rashba-Zeeman Hamiltonian for a chain of spinful single-level QDs can be expressed as follows
$H={}$ $\sum_{n,s,s^{\prime}}\Bigl[(-\mu_{n}\sigma_{0}+V_{\mathrm{Z}}\sigma_{z})_{ss^{\prime}}\,c_{n,s}^{\dagger}c_{n,s^{\prime}}\Bigr$.
$\left.+\tfrac{1}{2}\left(\left(\Delta_{n}e^{i\phi_{n}}i\sigma_{y}\right)_{ss^{\prime}}c_{n,s}^{\dagger}c_{n,s^{\prime}}^{\dagger}+\mathrm{h.c.}\right)\right$.
$+\Bigl.\left(t_{n}(e^{i\bm{\lambda}_{n}\cdot\bm{\sigma}})_{ss^{\prime}}c_{n,s}^{\dagger}c_{n+1,s^{\prime}}+\mathrm{h.c.}\right)\Bigr],$ (1)
with the on-site potentials $\mu_{n}$, inter-dot hopping amplitudes $t_{n}$ modulated by (Rashba) spin-orbit vector $\bm{\lambda_{n}}=\lambda_{n}(\sin\rho_{n}\cos\xi_{n},\sin\rho_{n}\sin\xi_{n},\cos\rho_{n})$, Zeeman energy $V_{\mathrm{Z}}$, proximity-induced superconducting ($s$-wave) pairing $\Delta_{n}e^{i\phi_{n}}$, and $n$ describing dot number while $s$ being spin degree of freedom. The values of those parameters are then carefully adjusted to have MZMs emerge in the system. One set of tuned (yet uniform) parameters (which we call reference parameters, $P_{0}$) for $N=3$ QDs chain includes: $\mu=0.6\,\mathrm{meV}$, $t=0.25\,\mathrm{meV}$, $\lambda=0.27\,\pi$, $\rho=\xi=\frac{\pi}{2}$, $V_{\mathrm{Z}}=0.5\,\mathrm{meV}$, and $\Delta=0.25\,\mathrm{meV}$. The Rashba length $\lambda=0.27\,\pi$ was tuned so that at $\mu=0.6$ meV two energy levels touch at zero energy (c.f. Fig. 5(a)). Also note that some parameters, i.e., $\{\mu_{n}$, $t_{n}$, $\lambda_{n}\}$ (7 in total) are easier to tune – electrically (via local gating) than the others: the global Zeeman field $V_{\mathrm{Z}}$ and the proximity-induced superconducting gap $\{\Delta_{n}\}$.
To train the ViT model in physics-informed manner, we introduce a differentiable quasi-metric $\mathcal{M}$, quantifying how close a given Hamiltonian is to MZM regime. The proposed phenomenology combines several physical indicators: edge-state localization, zero-energy spectral weight, and parity (electron-hole) symmetry. For a detailed $\mathcal{M}$ definition, see the Methods section. We also note that we tested the popular Majorana polarization [Sticlet2012, Maska2017, Szumniak2017] measure, directly related to the parity operator, but it yielded unsatisfactory training results. This is because it fails to discriminate topologically trivial zero modes that localize in the central QD or anti-localize in the left and right QDs (see the Supplementary Information of Ref. [Bordin2025] for examples of such states). We define our measure in a way that also discriminates these states.
We start with random set of parameters $P$ and collect conductance maps $G(H(P))$ – details on $G$ calculation can be found in the Methods section. PINNAT fed by $G$ maps predicts corrective updates $\delta P$ to a subset of Hamiltonian parameters. We assume that, depending on the experimental setup, different subsets of parameters may be available for tuning. Two versions of the PINNAT model to mimic different experimental arrangements was trained. The first one is trained to predict corrections to electrically controlled local parameters $\{\mu_{n},t_{n},\lambda_{n}\}$ – each of which can be adjusted independently, while leaving $\{\Delta_{n},V_{\mathrm{Z}}\}$ as background variables. The second one is allowed to adjust $\{\mu_{n},V_{\mathrm{Z}}\}$ pair, where $\mu_{n}$ can be corrected locally, while $V_{\mathrm{Z}}$ is adjusted globally. We also have to keep in mind that PINNAT model (as well as the experimentalist) does not know which parameters were detuned.
Maximizing $\mathcal{M}$ drives the NN to adjust $P$ such that edge states become increasingly localized and the strong mid-gap energy signal emerges, which is visualized in Fig. 1 by output conductance maps (this example is further analyzed in Fig. 5(d)). Specifically, PINNAT is trained to minimize the loss:
$\mathcal{L}(H(P^{\prime}))=\alpha\left\langle\delta P\right\rangle^{2}-\mathcal{M}(H(P^{\prime})),$ (2)
for the Hamiltonian with tuned parameters $H(P^{\prime})=H(P\,+\,\delta P)$. Additionally, to force the PINNAT to predict smallest possible correction $\delta P$ to the parameters $P$, an extra regularization term $\alpha\langle\delta P\rangle^{2}$ is included, with a factor $\alpha=0.1$. Training details, including parameters sampling used to generate synthetic training set of conductance maps, are described in Methods section.
III Results
Figure 2: $\mathcal{M}$ metric map for uniform (across the dots) change of different parameters – resulting in shifted $P$ – for (b) model adjusting $\{\mu_{n},t_{n},\lambda_{n}\}$, and (c) adjusting $\{\mu_{n},V_{\mathrm{Z}}\}$. The ranges of the parameter sampling that were used in the model training set are marked with a gray rectangle. The reference parameter values $P_{0}$ are marked with a red dot. In (a), we present values of $\mathcal{M}(H(P))$ before parameter tuning, while in (b) and (c) $\mathcal{M}(H(P+\delta P))$ – after tuning.
To present the PINNAT performance in parameter tuning, we plot the values of $\mathcal{M}$, disturbing selected three-QDs Hamiltonian parameters. Specifically, we shift (detune) the value of the given parameters (the rest is kept default, $P_{0}$) – obtaining $P$ – and verify what corrections $\delta P$ are predicted by the NN models. Then the $\mathcal{M}(H(P+\delta P))$ value is plotted. In Fig. 2(b) and (c) we show how $\mathcal{M}$ changes with uniform (i.e., the same for each QD) parameter shift for models tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ and $\{\mu_{n},V_{\mathrm{Z}}\}$, respectively. The values for $\mathcal{M}(H(P))$, i.e. before correction, are presented as reference in column (a). For example, in $t$ vs $\mu$ plot we shift parameters to $t_{1}=t_{2}=t$ and $\mu_{1}=\mu_{2}=\mu_{3}=\mu$ obtaining $P$ (column (a)), then predict correction $\delta P$ and present $\mathcal{M}(H(P+\delta P))$ in respective positions in columns (b) and (c). Additionally, in column Fig. 2(a), orange dashed curves mark the analytical conditions: (1) $\Delta=\sqrt{V^{2}_{\mathrm{Z}}-\mu^{2}}$ for the existence of two zero modes in decoupled QDs, and (2) $\lambda=\arctan(\frac{\mu}{\Delta})$ corresponding to the KC sweet spot (see the Supplementary Notes (SN) for details).
Results in In Fig. 2(b) and (c) show that the PINNAT models can effectively learn to identify various $H(P)$ regimes from $G$ maps and use it to significantly increase – in comparison to Fig. 2(a) – the regions with $\mathcal{M}>0$ by proposing corrections $\delta P$ to the subset of the parameters. Unsurprisingly, the models are most effective in the parameter regime covered in the training data, although some ability to generalize on unseen parameter ranges can be noticed, especially for $t$ and $V_{\mathrm{Z}}$ shift in Fig. 2(b). On the other hand, the PINNAT that updates the $\{\mu_{n},t_{n},\lambda_{n}\}$ exhibits difficulties in some regions in predicting appropriate corrections for the values of $t$ and $\lambda$, even though they are allowed to be modified.
Interestingly, the model that adjusts $\{\mu_{n},V_{\mathrm{Z}}\}$ – Fig. 2(c) is slightly more effective in proposing corrections, both in terms of parameter range that can be effectively corrected and the probability of MZMs emergence, as indicated by higher values of $\mathcal{M}$. This behavior is specifically worth highlighting as the model does not modify the $t_{n}$, $\lambda_{n}$ and $\Delta$ parameters, while still possessing the ability to effectively amend them by updating $\mu_{n}$ and $V_{\mathrm{Z}}$. These results show that $V_{\mathrm{Z}}$ plays a more important role in controlling the system than the hoppings $t_{n}$ and SOI $\lambda_{n}$. This is consistent with observations known from long chains [stanescu2024, Leumer2021], where the magnetic field is essential to compete with the superconducting pairing to drive the system into the topological phase. However, it should be noted that when the hopping parameters $t$, $\lambda$ are close to zero in the first and second rows of Fig. 2(c) the model likewise selects symmetric zero modes localized at the edge QDs with a high $\mathcal{M}$; nevertheless in this case (in the limit $t\rightarrow 0$) the system is clearly in a trivial phase [stanescu2024].
Figure 3: Same as in Fig. 2: $\mathcal{M}$ metric maps (a) before, and (b,c) after parameters tuning, but now for local shift of selected pair of parameters.
PINNAT behavior is further investigated in Fig. 3 showing how $\mathcal{M}$ changes with (a pair of selected) local parameters shift. In case of a shift in the $\mu_{n}$ parameter, both models produce comparable results – although tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ allows one to cover a wider range of parameters, adjusting $\{\mu_{n},V_{\mathrm{Z}}\}$ allows to reach higher values of $\mathcal{M}$. Simultaneously, sweeping $t_{n}$ and $\lambda_{n}$ produces larger discrepancies in favor of the PINNAT model that corrects the parameters $\{\mu_{n},V_{\mathrm{Z}}\}$. Nevertheless, both models demonstrate the ability to effectively correct the dot-specific parameters and increase the probability of observing desired MZM modes.
Figure 4: Iterative autotuning procedure for (a) model adjusting $\{\mu_{n},t_{n},\lambda_{n}\}$, and (b) adjusting $\{\mu_{n},V_{\mathrm{Z}}\}$. In the first (left) plot, we present $\mathcal{M}$ map before tuning. Then, in the center, parameters are tuned with a single step of NN corrections. Finally, in the last (right) plot, there is a map plotted for parameters tuned within 10 subsequent steps of NN corrections.
Since the models tend to work in a rather limited area of parameters covered in the training data, we propose an iterative modification to the proposed procedure. After applying the initial corrections proposed by PINNAT, one can measure the conductance maps for the corrected system and pass them to the NN in the subsequent step of autotuning. As a result, in Fig. 4, we can notice that within 10 steps of corrections, the $\mathcal{M}$ metric can be significantly increased even in the regions with an initially zero value of $\mathcal{M}$.
Figure 5: Three-QD chain Hamiltonian $H$ as a function of the offset $\mu$ added to local potentials: $\mu_{n}\rightarrow\mu_{n}+\mu-0.6$ meV (dashed vertical line marks the reference $\mu=0.6\,\mathrm{meV}$). Additionally $\mathcal{M}$ metric is plotted, and eigenvalues are colored (left column) with edge occupations – 1 means that the state is localized on the edges, while 0 means that state is localized in the center dot, or (right column) electron-hole symmetry – calculated as difference between density of electrons and holes. Plots are presented for (a) reference parameters; (b) modified $\lambda_{1}$ and $\lambda_{2}$; (c) parameters from (b) but with NN-tuned $\{\mu_{n},t_{n},\lambda_{n}\}$; (d) parameters from (b) with NN-tuned $\{\mu_{n},V_{\mathrm{Z}}\}$.
Finally, we present a specific example of the autotuning procedure for the $\lambda_{n}$ parameter. In Fig. 5, we subsequently plot the eigenvalue spectrum of the three-QD chain Hamiltonian $H$ as a function of the global offset $\mu$ added to local potentials $\mu_{n}+\mu$. In Fig. 5(a) we have exactly reference parameters $H(P_{0})$ case, where we can clearly observe MZMs separated (by a topological gap) from the rest of the spectrum with maintained electron-hole symmetry and (partially) localized on the edges. Conductance maps for the reference parameters, $P_{0}$ case: $G(H(P_{0}))$ are presented in Fig. 6.
Adding a noise to $\lambda_{n}$ (by setting $\lambda_{0}=0.62\,\pi$ and $\lambda_{1}=0.41\,\pi$) – presented in Fig. 5(b) – results in reduced $\mathcal{M}$ value and vanishing MZMs. Then the noisy parameter can be corrected using either of the two presented NN models. Tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ – presented in Fig. 5(c) – makes MZMs reappear with desired properties, i.e., they are localized on the edges, electron-hole symmetry is preserved, and the edge states are slightly gapped from the rest of the spectrum. On the other hand, tuning $\{\mu_{n},V_{\mathrm{Z}}\}$ – Fig. 5(d) – not only leads to restoring MZMs at the exact point in the tuned parameter space with the larger topological gap but also increases the probability of observing MZMs in a wider range of $\mu$ values. An additional example of analogous autotuning, but for the $t$ parameter is provided in SN.
IV Discussion
These findings align with and extend current efforts in ML-assisted control of Majorana platforms [Thamm2024, Taylor2025disordermitigation, vandriel2024]. Most notably, the recent work of Taylor and Das Sarma, Ref. [Taylor2025disordermitigation] introduced a ViT-based framework for mitigating strong disorder in Majorana nanowires. Their study demonstrated that disorder can be compensated – sometimes restoring topology in systems that originally show no nontrivial regions – using conductance-based NN inference combined with evolutionary optimizer (CMA-ES). Our approach differs in several important respects. First, instead of relying on a second-stage optimizer such as CMA-ES, our network learns Hamiltonian behaviors manifested by transport measurements to directly predict parameter corrections, making the tuning process faster and conceptually closer to a experimental tuning protocol. Second, whereas Ref. [Taylor2025disordermitigation] focuses primarily on restoring scattering-invariant and LDOS-based indicators in bulk nanowires, our work addresses QD chains, a complementary Majorana platform with discrete QDs and gate-defined tunnel couplings between them. Third, our method integrates NN physics-informed via differentiable quasi-metric $\mathcal{M}$ capturing edge localization, electron-hole symmetry, and spectral proximity to zero energy; this contrasts with supervised learning to pretict topology-based indicators.
We also emphasize that our results highlight the possibility of generalization beyond the training regime. For several parameters, notably $t$ and $V_{\mathrm{Z}}$, the networks propose meaningful corrections even when initial values fall outside the training set distribution. Specifically, after increasing the number of consecutive steps of corrections, the network is capable of effective autotuning in a broad area of parameters. This indicates that unsupervised PINNAT architecture is capable of learning physically sensible parameter correlations rather than just approximating between samples (Hamiltonians) from the training set. Such generalization is essential if these methods are to be deployed on real devices, where fabrication imperfections and drift may move operating points far from any simulated training set.
Importantly, the proposed PINNAT scheme is fully physics-informed: MZM metric $\mathcal{M}$, together with the the Hamiltonian family definition and eigensolver implementation, is explicitly encoded in the structure of the NN loss function. This approach aligns with the recent strong trend toward designing NNs that respect the underlying physics of the problem [Bronstein2021], for example in studies of quantum entanglement [Pawlowski2024, Krawczyk2024], quantum control [Norambuena2024] or in the optimization of quantum tomography protocols [Krawczyk2025].
While our current proof-of-concept study is restricted mainly to a three-QDs system, we show the ability to scale the proposed method to 7-QDs chain in the SN. With some minor adjustments in the training procedure we claim that it is possible to have the PINNAT model trained to propose significant corrections in the larger systems, maintaining its ability to autotune the Hamiltonian parameters beyond the training regime.
The major intrinsic limitation of this work is that our method relies on simulated conductance maps, whereas experimental data may include additional noise sources, nonidealities, and systematic offsets not captured in the numerical model. Future work should therefore incorporate synthetic noise or hybrid train-on-simulation–fine-tune-on-experiment protocols to bridge this gap.
V Conclusions
In this work, we demonstrated that ViT-based neural network physics-informed by $\mathcal{M}$ metric, trained on conductance maps can effectively autotune QDs chain Hamiltonian toward the emergence of MZMs. Our results show that the proposed framework succeeds in correcting both global deviations of system parameters – those that shift all QDs uniformly – and local parameter noise that affects individual QDs independently. In both regimes, PINNAT consistently increases the Majorana metric $\mathcal{M}$, thereby restoring the formation of zero-energy edge states where possible.
Proposed framework merges quantum transport simulation with ML-based parameter feedback for topological phase tuning. By integrating experimental observables (conductance maps) as model input and theoretical descriptors ($\mathcal{M}$ measure) as physics-informed targets for model training, our approach paves the way for autonomous tuning into robust topological regimes in noisy mesoscopic systems.
VI Methods
Proposed method uses conductance maps as high-dimensional visual inputs for the ViT model. The conductance $G$ is calculated using the $S$-matrix formalism in the wide-band limit [Donarini2024] via the Weidenmüller formula [Bordin2025, christiansen2009] for $N=3$ QDs chain:
$S(E_{F})=\begin{pmatrix}s^{ee}&s^{eh}\\ s^{he}&s^{hh}\end{pmatrix}=\mathbf{1}-iW^{\dagger}(E_{F}-H+\frac{i}{2}WW^{\dagger})^{-1}W,$ (3)
with the tunneling matrix defined as $W=\mathrm{diag}(1,0,1)\otimes\mathrm{diag}(\sqrt{\Gamma},\sqrt{\Gamma},-\sqrt{\Gamma},-\sqrt{\Gamma})$ and the dot-lead coupling $\Gamma=0.1$ meV. If we reshape the $S$-matrix accordingly $S=S_{n,p,s,n^{\prime}\!,p^{\prime}\!,s^{\prime}}$ ($n=L,C,R$ indexing dots, $p=1,2$ particle, and $s$ – spin subspaces) then the respective reflection matrices: $s^{ee}_{s,s^{\prime}}\!(i,j)=S_{i,1,s,j,1,s^{\prime}}$ and $s^{he}_{s,s^{\prime}}\!(i,j)=S_{i,2,s,j,1,s^{\prime}}$ give differential conductance as, $ss^{\dagger}\equiv|s|^{2}$,
$G_{ij}(E_{F})\equiv\frac{\mathrm{d}I_{i}}{\mathrm{d}V_{j}}=2\delta_{ij}-\mathrm{tr}\!\left(|s^{ee}(i,j)|^{2}\right)+\mathrm{tr}\!\left(|s^{he}(i,j)|^{2}\right)$ (4)
in unit of $e^{2}\!/h$, $i,j=L,R$ denoting left ($L$) or right ($R$) lead, and $E_{F}$ being the Fermi energy in the leads. The input maps include 4 conductance $G_{ij}$ components: $G_{\mathrm{LL}}$, $G_{\mathrm{LR}}$, $G_{\mathrm{RL}}$, and $G_{\mathrm{RR}}$, where for instance $G_{\mathrm{LL}}=\frac{dI_{\mathrm{L}}}{dV_{\mathrm{L}}}$, with $I_{\mathrm{L}}$ denoting current through the left lead, and $V_{\mathrm{L}}$ is the bias voltage of the left lead. Similarly, other components can be defined by using different combinations of left (L) and right (R) leads. Noteworthy, each conductance map is defined by 2D plot of $G_{ij}$ as a function of some parameter and $E_{F}$. We utilize 4 maps per each component: 3 for $\mu_{i=\mathrm{L},\mathrm{C},\mathrm{R}}$ variation and 1 for $V_{\mathrm{Z}}$ variation: 16 maps in total serving as the input. Fig. 6 shows the conductance maps ($G_{LL}$ component) for the reference parameters $P_{0}$, highlighting the emerged zero-biased peaks.
Figure 6: Conductance maps for the reference parameters $P_{0}$. Zero-bias peaks are indicated by orange arrows.
The $\mathcal{M}$ metric, utilized to train the PINNAT model, is the following:
$\mathcal{M}(H(P))$ $=\frac{p_{0}}{2}\max\!\left[0,\,2\,m_{0}-\sum_{i>1}m_{i}\right],$ (5)
$m_{i}$ $=|\langle\psi_{i}|M\rangle|e^{-|E_{i}|/\epsilon},$
$p_{i}$ $=2\max\!\left[0,\,4|u_{i}|^{2}|v_{i}|^{2}-\tfrac{1}{2}\right],$
with $\{\psi_{i},E_{i}\}$ being the collection of $H(P)$ eigenpairs. Both $m_{i}$ and $p_{i}$ are sorted in ascending order by $|E_{i}|$. Eigenstates projection $m_{i}$ on the left (or equivalently right) Majorana mode $M=\gamma_{L}$ is weighted by their distance from zero energy $|E_{i}|$ with some threshold $\epsilon=0.1$ meV. Second factor – $p_{i}$ quantifies electron-hole symmetry: $u_{i},v_{i}$ are the electron-hole components of eigenvectors $\psi_{i}$. By the $H$ symmetry $m_{0}=m_{1}$, and also $p_{0}=p_{1}$. For ideal MZM pair we should have $m_{0}=m_{1}=1$ and $p_{0}=p_{1}=1$, giving $\mathcal{M}=1$. The form of the metric, Eq. 5, is designed to discriminate trivial zero modes described e.g. in the SI of Ref. [Bordin2025], especially $\sum_{i>1}m_{i}$ term penalizes the presence of trivial zero modes, even when the topological ones are there as well.
In our work we adapt typical ViT architecture [dosovitskiy2021] adjusted to process 16-channel $50\times 50$ input tensor (representing all generated conductance map components) and returning vector of the Hamiltonian parameter corrections $\delta P$. The architecture hyperparameters such as number of attention heads (4), number of attention blocks (6), hidden size (256) and patch size (1) are tuned to decrease value of $\mathcal{L}$ after 100 epochs below given threshold $\tau=-0.2$.
During single epoch of training we iterate through 10 00010\,000 independent $H$ samples generated using Eq. 1, varying parameters $\{\mu_{n},t_{n},V_{\mathrm{Z}},\Delta_{i}\}$ in range $[0,1]$ $\mathrm{meV}$ and $\lambda_{n}$ in range $0,\pi]$. The training is proceeded until convergence of $\mathcal{L}$, Eq. 2, is observed – usually between 150 and 200 epochs.
Acknowledgements
We would like to thank M. Margańska-Łyżniak for interesting discussions and helpful comments on the manuscript. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2025/018433.